# Scaling Reproducibility: An AI-Assisted Workflow

### v1

@@ Abstract
Reproducibility is central to research credibility, yet large-scale reanalysis of empricial data remains costly because replication packages vary widely in structure, software environment, and documentation. We develop and evaluate an agentic AI workflow that addresses this execution bottleneck while preserving scientific rigor. The system separates scientific reasoning from computational execution: researchers design fixed diagnostic templates, and the workflow automates the acquisition, harmonization, and execution of replication materials using pre-specified, version-controlled code. A structured knowledge layer records resolved failure patterns, enabling adaptation across heterogeneous studies while keeping each pipeline version transparent and stable. We evaluate this workflow on 92 instrumental variable (IV) studies, including 67 with manually verified reproducible 2SLS estimates and 25 newly published IV studies under identical criteria. For each paper, we analyze up to three two-stage least squares (2SLS) specifications, totaling 215. Across the 92 papers, the system achieves 87% end-to-end success overall. Conditional on accessible data and code, reproducibility is 100% at both the paper and specification levels. The framework substantially lowers the cost of executing established empirical protocols and can be adapted in empirical settings where analytic templates and norms of transparency are well established.
Keywords: reproducibility, replication, research transparency, open science, AI-assisted workflows, agentic AI, Claude Skills, causal inference
@@ 1. Introduction
Reproducibility is fundamental to research credibility and cumulative scientific progress. In empirical social science, reproducible analyses allow researchers to verify published claims, scrutinize identifying assumptions, and assess the practical relevance of new methodological developments. As empirical methods evolve rapidly, access to real-world data and code has become increasingly important not only for assessing research credibility, but also for advancing methodology through systematic reanalysis of existing studies.
@@
Institutional norms have expanded the availability of replication materials. Leading journals in economics and political science now require authors to post data and code, and some conduct in-house replication checks before publication. Yet availability alone does not ensure reproducibility at scale. Replication packages vary widely in software environment, directory structure, naming conventions, documentation quality, and execution logic. Even when materials are public, reproducing results across many papers remains costly and fragile. The bottleneck is operational: executing idiosyncratic replication materials in a standardized and auditable manner requires substantial researcher time.
This paper develops and evaluates an agentic AI workflow to address this execution bottleneck. The workflow combines adaptive coordination with deterministic computation. A large language model (LLM) routes tasks across modular agents that ingest replication materials, identify specifications, reconstruct computational environments, execute models, and generate standardized diagnostic reports. A structured knowledge layer records previously resolved failure patterns and clarifies stage-level responsibilities, allowing the system to accumulate experience across studies while keeping each pipeline version transparent and stable. All numerical operations—data preparation, estimation, and diagnostic computation—are carried out by version-controlled program code. For a fixed pipeline version and fixed inputs, reruns produce identical numerical outputs and retain a complete audit trail of intermediate artifacts and logs. The paper does not propose new estimators or diagnostics. Instead, we ask whether established empirical protocols can be executed reliably and at scale under real-world conditions. Our evidence suggests that they can.
A central design principle of this workflow is the separation of scientific reasoning from computational execution. Human researchers design diagnostic templates that specify estimands, estimators, robustness checks, and summary measures appropriate for a given research design. Once these templates are fixed, reproduction largely consists of execution-oriented tasks: acquiring replication packages, reconstructing computational environments, locating and running prespecified specifications, extracting analysis datasets, and harmonizing outputs. At the current stage of development, AI systems can not design diagnostic tools that meet the precision standards implied by econometric and statistical theory. We therefore treat diagnostics as human-designed inputs and evaluate whether AI can execute them reliably and reproducibly at scale. This division of labor may evolve as AI systems improve, but it aligns with current research needs.
We evaluate the workflow on a corpus of 92 studies with instrumental variable (IV) designs. Of these, 67 were previously analyzed in Lal et al. (2024), where the authors manually verified the reproducibility of at least one the two-stage least squares (2SLS) coefficient in each study. We extend the analysis to 25 additional IV studies published after the original sample, applying identical inclusion criteria and the same diagnostic template. Across the combined corpus, the workflow targets up to three 2SLS specifications per paper. Each specification corresponds to a model defined by an outcome, a single treatment variable, one or more instruments, and a set of covariates, estimated on a particular sample. In the expanded set of 92 studies, the system achieves a 87% end-to-end reproducibility success. The unsuccessful cases are caused by incomplete replication materials rather than computational instability. Conditional on accessible materials, the pipeline reproduces the benchmark 2SLS estimates exactly and completes all the diagnostic tests.
@@
It is important to note that this level of reliability was not achieved in a single engineering pass. The corpus spans multiple programming languages (mostly Stata and R), estimation commands, directory structures, and idiosyncratic coding practices. Many failure modes arise only when new replication packages are encountered. We therefore adopt an adaptive, human-in-the-loop process. When a recurring failure pattern is identified, it is encoded as a generalized rule in the execution layer and version-controlled between runs. Coverage expands across versions, while numerical behavior remains fixed within each version. The appendix documents the classes of variability encountered and the corresponding adjustments that resolved them.
This adaptive process reflects the challenges we encountered in our prior large-scale reanalysis projects. In earlier work, we manually reconstructed and reanalyzed dozens of published studies using interaction models, IV designs, and parallel trends designs using panel data (Hainmueller et al., 2019; Lal et al., 2024; Chiu et al., 2023). Each project required years of coordinated effort. Much of that time was devoted not to methodological development, but to deciphering authors’ codebases, repairing path dependencies, reconstructing environments, and harmonizing heterogeneous replication materials. Those experiences revealed that once a diagnostic template is specified, most remaining work is procedural—and therefore, in principle, automatable. They also produced structured benchmark corpora that enable systematic evaluation of an automated executor.
Viewed more broadly, this project connects to long-standing concerns about research credibility. Calls to improve empirical practice—ranging from Leamer’s appeal to “take the con out of econometrics” (Leamer, 1983) to the credibility revolution in economics (Angrist and Pischke, 2010) and political science (Torreblanca et al., 2026)—have emphasized transparency, auditability, and disciplined empirical workflows alongside advances in causal identification. More recent discussions of the replication crisis, particularly in psychology (Open Science Collaboration, 2015), highlight the consequences of fragile research pipelines. In this paper, we adopt the now-standard distinction between reproducibility—recovering reported results using the original data and code—and replicability—obtaining similar findings in new studies. Our focus is explicitly on reproducibility. It is not a substitute for replication, but a necessary first step toward credible inference and cumulative research.
The contributions of this paper are threefold. First, we design and implement an adaptive yet version-controlled agentic AI workflow that executes fixed empirical templates across heterogeneous replication materials. Second, we provide systematic evaluation against a manually verified benchmark and a forward extension to newly published studies, documenting a 100% end-to-end success rate in the expanded corpus. Third, we make transparent the variability inherent in real-world replication packages and the engineering adjustments required to accommodate it, offering a disciplined template for similar efforts in other research designs.
@@
The scope of this study is currently limited to empirical social science, particularly causal inference, with data drawn primarily from political science, where norms for sharing replication materials are relatively strong. The workflow depends on the availability of usable code and data and is not designed to recover results when replication materials are missing, incomplete, or fundamentally flawed. The findings therefore speak to what is feasible under current best practices rather than to settings without accessible materials. At the same time, the framework itself is not discipline-specific. Wherever empirical fields have established analytic templates and norms of transparency, the same template–executor approach can be adapted to support large-scale reproducibility and systematic reanalysis.

### v2

@@ Abstract
Computational reproducibility is central to scientific credibility, yet verifying published results at scale remains costly. We develop an AI-assisted workflow for automated full-paper replication—retrieving materials, reconstructing environments, executing code, and matching outputs to point estimates reported in regression tables. We define a universe of all empirical and quantitative papers from the three top political science journals (2010–2025) and measure stated data availability using automated extraction. For a stratified sample of 384 studies, we apply the workflow to conduct full-paper replication, totaling 3,382 empirical models. We find that journal verification requirements, combined with data archiving mandates, drive reproducibility: the full-paper reproducibility rate rises from 29.6% before DA-RT adoption to 79.8% after, and conditional on accessible replication packages, 94.4% of papers are fully reproducible (237/251). As a secondary application, we apply standardized IV diagnostics to 92 studies (215 specifications), illustrating how automated execution enables systematic reanalysis across heterogeneous empirical settings.
Keywords: reproducibility, replication, research transparency, open science, AI-assisted workflows, agentic AI, causal inference
@@ 1. Introduction
Reproducibility is fundamental to research credibility and cumulative scientific progress. In empirical social science, reproducible analyses allow researchers to verify published claims, scrutinize identifying assumptions, and assess the practical relevance of new methodological developments. Access to real-world data and code has therefore become important both for assessing research credibility and for advancing methodology through systematic reanalysis.
@@
Institutional norms have expanded the availability of replication packages—bundles of data, code, and documentation deposited by authors to enable reproduction of published results. Leading journals in political science and economics now require authors to post replication packages, and some conduct replication checks before publication. Yet availability alone does not ensure reproducibility at scale. Replication packages vary widely in environments, structure, and execution logic, so reproducing results across many papers remains costly and fragile. The bottleneck is operational: executing heterogeneous packages in a standardized and auditable manner requires substantial effort.
This paper develops an AI-assisted workflow for automated full-paper replication to address this execution bottleneck. The workflow reconstructs computational environments, executes code, and verifies published results. Its purpose is to determine whether existing empirical analyses can be reproduced reliably and at scale, not to introduce new estimators or diagnostics. Once data and code are standardized through this process, they can be readily reused, lowering the cost for methodological researchers to develop and apply new methods on a common empirical basis.
A central design principle is the separation of scientific reasoning from computational execution. Researchers specify the empirical targets to be reproduced—model specifications and reported quantities of interest as defined by the paper’s narrative and replication package. Conditional on these inputs, replication reduces to execution-oriented tasks: acquiring replication packages, reconstructing environments, locating and running specifications, extracting analysis datasets, and harmonizing outputs. Throughout this paper, we use reproducibility to mean re-execution of existing code and data (National Academies of Sciences, Engineering, and Medicine, 2019), distinguishing it from replicability—obtaining similar findings in new studies. Earlier literature sometimes uses “replication” for the former (King, 1995). [^1] The terminology varies across fields. In political science, King (1995) used “replication” to encompass both re-execution of existing analyses and new studies testing the same hypothesis. Economics has a similar tradition (Vilhuber, 2020). We follow the distinction codified by the National Academies (National Academies of Sciences, Engineering, and Medicine, 2019) and adopted by recent work in both economics (Vilhuber, 2020) and political science (Alvarez and Heuberger, 2022): reproducibility denotes re-execution of the original code and data, while replicability denotes obtaining consistent findings with new data. The compound “replication package”—the standard term for the bundle of code and data deposited by authors—retains the older usage; the property we measure when re-executing those packages is reproducibility.
Using this workflow, we study a large set of top political science publications. We define the universe as all empirical and quantitative papers in the American Political Science Review (APSR), American Journal of Political Science (AJPS), and Journal of Politics (JOP) from 2010 to 2025, totaling 3,464 papers. For this universe, we measure stated data availability using automated extraction of availability statements. Combined with replication outcomes, this allows us to compare trends in stated availability and actual reproducibility over time and relate them to changes in journal policies, particularly the introduction of verification requirements.
We then apply the workflow to two tasks. First, we conduct full-paper replication for a stratified sample of 384 studies (8 papers per journal-year). For each paper, we execute the complete codebase and attempt to reproduce point estimates reported in regression tables, excluding figures, descriptive tables, and simulation results. This defines a consistent unit of evaluation across heterogeneous designs and software environments. The estimated full-paper reproducibility rate rises from 29.6% before DA-RT adoption to 79.8% after, using journal-specific implementation dates (APSR: 2016; AJPS and JOP: 2015). This increase is driven by journal policies requiring public data archiving and, especially, in-house or third-party verification. Conditional on accessible materials, 237 of 251 papers (94.4%) are fully reproducible and 239 of 251 (95.2%) are fully or largely reproducible regardless of publication year—failures trace almost entirely to missing or restricted data, not to code quality.
Second, we apply the workflow to instrumental variable (IV) designs. Building on Lal et al. (2024), we analyze 92 studies and extract 215 specifications for standardized diagnostics, illustrating how automated execution enables systematic analysis across heterogeneous empirical settings. Among the 215 specifications, the workflow achieves an 82% end-to-end autonomous success rate (55 of 67 original papers); all failures reflect unavailable replication archives rather than execution errors. Conditional on accessible materials, the workflow reproduces benchmark estimates exactly and completes all diagnostic tests.
@@
This level of reliability is achieved through an adaptive, version-controlled process. Replication packages span multiple programming languages, estimation commands, directory structures, and coding conventions, and many failure modes arise only when new materials are encountered. Recurring issues are encoded as generalized rules in the execution layer and versioned across runs. Coverage expands across versions, while numerical behavior remains fixed within each version. The Supplementary Materials document these failure patterns and the corresponding adjustments.
In political science, King (1995) argued that scholarly norms should require sufficient data and code to reproduce published results. This vision was formalized through the Data Access and Research Transparency (DA-RT) initiative and the 2014 Journal Editors’ Transparency Statement (JETS), building on public data repositories—principally the Harvard Dataverse (King, 2007), ICPSR, and the Open Science Framework—that provided the infrastructure for deposit and retrieval. The three journals in our corpus adopted verification requirements at different times (Data Access and Research Transparency (DA-RT), 2014; Key, 2016), providing institutional variation that allows us to assess whether these policies are associated with improved reproducibility in practice.
These transparency norms made possible a series of large-scale reanalysis projects (Hainmueller et al., 2019; Lal et al., 2024; Chiu et al., 2026), in which much of the effort was devoted to reconstructing environments and harmonizing replication packages rather than developing new methods. Those experiences suggest that, once empirical targets are specified, most remaining work is procedural and therefore automatable. The present workflow operationalizes this insight.
More broadly, this project connects to long-standing concerns about research credibility, from Leamer’s call to “take the con out of econometrics” (Leamer, 1983) to the credibility revolution (Angrist and Pischke, 2010; Torreblanca et al., 2026) and the replication crisis in psychology (Open Science Collaboration, 2015). These discussions emphasize verifiable analysis and expose the consequences of fragile research pipelines.
The main contribution of this paper is twofold. First, we make computational reproducibility measurable at scale and document that journal transparency policies—particularly verification requirements—are associated with substantially higher reproducibility rates in political science. Second, we show that automated execution enables standardized diagnostics across published studies, illustrated with IV designs, and can facilitate methodological research by providing reusable data and outputs.
@@
The scope of this study is empirical social science, with data drawn from top political science journals where norms for sharing replication packages are relatively strong. The workflow requires usable code and data and does not recover results when packages are missing or fundamentally flawed. The findings therefore reflect what is feasible under current best practices. Although the analysis focuses on political science, the framework is general and can be extended to other fields with established transparency norms.
2. AI-Assisted Reproducibility Workflow
This section describes the AI-assisted workflow used to conduct full-paper replication and design-specific diagnostics. The workflow targets the execution bottleneck identified in the introduction: executing heterogeneous replication packages in a standardized and auditable manner. It does not automate methodological reasoning or introduce new statistical procedures. Instead, it standardizes and accelerates execution when usable data and code are available.
Design principles.
Large-scale reproducibility involves a basic tension between heterogeneity and determinacy. Replication materials vary widely across studies in programming language (Stata, R, and Python), directory structure, naming conventions, and documentation quality. At the same time, reproducibility requires determinacy: for a fixed pipeline version and fixed inputs, numerical outputs must not depend on ad hoc decisions, platform-specific defaults, or stochastic behavior.
The workflow resolves this tension through three separation principles. First, adaptive coordination is separated from fixed computation. An LLM orchestrator routes tasks, interprets failures, and selects among predefined recovery steps. All numerical work—data preparation, model estimation, and diagnostic computation—is executed by version-controlled program code. For a fixed pipeline version and fixed inputs, the workflow produces identical numerical outputs and retains a complete audit trail. Second, scientific reasoning is separated from execution. Researchers define the empirical targets and, where applicable, the diagnostic templates. The workflow handles the operational tasks: acquiring replication packages, reconstructing environments, executing code, and harmonizing outputs. Third, execution is separated from verification. Running an author’s code successfully does not by itself establish that the paper’s reported results have been reproduced. The workflow therefore treats code execution and result verification as distinct phases, with an explicit gate between them.
Architecture.
The system is organized as a three-layer architecture. At the top layer, a large language model (LLM), Claude, serves as an orchestrator that dispatches tasks, interprets errors, and determines how the pipeline proceeds. The middle layer consists of structured skill descriptions that define each stage’s input–output contract and record previously resolved failure patterns. At the bottom layer, deterministic agent code and statistical scripts perform all file operations and computation. The orchestrator controls task routing but is excluded from numerical estimation and inference. Figure LABEL:fig:architecture illustrates the architecture.
Execution proceeds through three phases. In Phase A (acquisition and execution), the system extracts metadata and replication links from the paper’s PDF, downloads the replication package from public repositories, prepares the code for automated execution by resolving path dependencies and environment assumptions, and executes the complete codebase. All estimation commands encountered during execution are instrumented to capture structured output, including coefficients, standard errors, sample sizes, and clustering. The output is a comprehensive log of every regression the code produces.
In Phase B (reproducibility verification), the system extracts all coefficients reported in the paper’s regression tables and compares them against the code-generated outputs. Matching uses precision-aware rounding: each code coefficient is rounded to the same number of decimal places as the paper-reported value, then checked for exact equality. Optimal assignment across paper–code coefficient pairs ensures that each reported value is matched to the most appropriate code output. Phase B concludes with a reproducibility verdict: papers classified as fully or largely reproducible proceed to Phase C; others enter a bounded fix loop (up to three iterations) before the verdict is finalized.
In Phase C (diagnostic evaluation), the system identifies the research design used in each reproduced specification—currently supporting IV and difference-in-differences designs—and applies the corresponding diagnostic template. For IV designs, this includes first-stage FF-statistics, Anderson–Rubin tests, bootstrap confidence intervals, jackknife sensitivity, and OLS comparisons, implemented through the ivDiag package (Lal and Xu, 2024). Phase C concludes with a verification check confirming that all diagnostics completed without error. The system then compiles a standardized report combining the reproducibility assessment from Phase B and the diagnostic results from Phase C.
Each phase produces explicit intermediate artifacts—structured logs, extracted datasets, matching reports, diagnostic outputs—written to disk. Agents share no hidden state. This modularity ensures that every step is inspectable, and execution can resume from any phase without restarting the pipeline.
Adaptation.
The workflow was developed iteratively through repeated encounters with diverse replication packages. Replication materials span multiple programming languages, estimation commands, directory structures, and coding conventions, and many failure modes arise only when new materials are encountered. When a recurring failure pattern is identified, it is encoded as a generalized rule in the execution layer and version-controlled between runs. Coverage expands across pipeline versions, while numerical behavior remains fixed within each version. The executor is kept design-agnostic: extending to new research designs requires specifying a new diagnostic template rather than modifying the underlying architecture.